Find both the circumference and the area of the circle whose equation is (x-2)^2 + (y-3)^2=64. Approximate your answer to the hearest hundreth of a unit.
1
Expert's answer
2012-06-05T09:39:06-0400
circumference = 2*pi*r
area=pi*r^2
where r - radius of the circle
from
equation radius is equal to sqrt(64)=8
so

circumference=2*pi*8=50.27
area=pi*64=201.06